An investment of 5600 is made at an annual simple interest rate of 6.5%. How much additional money must be invested at an annual simple interest rate of 9% so that the toatl interest earned is 8% of the total investement
Answer by checkley77(12569)

5600*.065+.09x=.08(5600+x)
364+.09x=448+.08x
.09x-.08x=448-364
.01x=84
x=84/.01
x=8,400 invested @ 9% will stiafy the need.
proof:
5600*.065+.09*8,400=.08(5600+8,400)
364+756=.08*14,000
1,120=1,120